# From Lab to Field: Real-World Evaluation of an AI-Driven Smart Video Solution to Enhance Community Safety

### v2

@@ Introduction
Record
Time
Camera Class ID
Bounding
Box
Feature Local ID Global ID
Anomaly
Score
@@
23:59:59 1 0 [x, y, w, h] Tensors $9$ 1001 40
Cloud Module
@@
leverages cloud-native services for data storage, management, and API generation (Dahunsi, Idogun, and Olawumi 2021). It minimizes lag between anomaly detection and notification, using a rule-based messaging service to send real-time alerts via email, text, or app notifications. A low-latency database supports real-time data access, while APIs generated through an application development kit optimize data retrieval (Mohammed et al. 2022; Ardabili et al. 2023b).
End User Devices
@@
is designed to promptly notify users of detected anomalies via a smartphone application (Ardabili et al. 2023a). The app provides real-time data and analysis, ensuring consistent functionality across different devices and operating systems, thereby enhancing accessibility and delivering information efficiently.
Figure 4: This represents an example of 4 pipelines running in the testbed with IP cameras deployed at different locations.
Deployment and Setup
@@
In lab settings, each advanced software features demonstrates significant potential individually; however, to fully realize its benefits, a comprehensive end-to-end evaluation in a real-world environment is essential. Such an evaluation allows for assessing the system’s load handling and endurance under actual conditions, ensuring that it can perform reliably and effectively when deployed, which might not be evident in a controlled environment.
In our study, an existing network of 16 AXIS IP cameras operating at 30 FPS with 720p resolution across the college campus is applied, covering approximately 35,000 square feet indoors and 60,000 square feet outdoors. Figure 3 visually represents the camera placements. Three cameras monitor outdoor areas, while the remaining 13 oversee various indoor locations, including entry points, vending machines, hallways, and communal spaces. The indoor cameras are mounted at 7 feet 6 inches, and the outdoor cameras at 10 feet 8 inches. Each camera has a varifocal lens with a horizontal field of view of 100 to 36 degrees. The system operates on a dedicated server equipped with a 16-core CPU, 252 GB of memory, and four GPUs, each with 24 GB VRAM.
@@
As shown in Figure 4, a sample output from the Local Node during a scenario with four operational cameras is presented, with privacy-preserving instance segmentation applied.
Applications and Visualizations
This section illustrates how the data collected from real world can be analyzed and visualized to enhance situational awareness in applications such as urban planning, resource allocation, and crowd management.
Descriptive Data
@@
In our descriptive analysis, the Global ID from human feature data serves as the unique identifier for tracking across data streams. We present five key metrics to provide a general overview of traffic flow and occupancy trends: real-time count of people, hourly average number of people per camera per location, total number of people and peak hour analysis over time. These metrics help to monitor occupancy, understand distribution trends, and identify peak traffic times, aiding in resource allocation and emergency planning.
Situational Awareness
@@
Situational awareness is crucial for effective decision-making across various domains. Our study explores four following key visualization techniques:
Occupancy Indicator
@@
provides insight into the number of people in a specific location by comparing current occupancy levels to historical data, which particularly useful in emergencies or public health scenarios. To calculate occupancy, the system compares current detections with historical percentiles, categorizing occupancy as ”Low,” ”Normal,” or ”High.”
Figure 5: Comparing different views of one camera. The graph (a) generated directly from database shows the original camera view. The graph (b) represents the average bird’s eye view coordinates.
(a) Camera 2- Weekday
(b) Camera 2- Weekend
(c) Camera 7- Weekday
(d) Camera 7- Weekend
Figure 6: Heatmap representation of the two cameras during weekday and weekends. Graph (a) and (c) represent the heatmaps during weekdays, while graph (b) and (d) show congestion of the same areas during weekends.
Statistical Anomaly
@@
often indicate unusual traffic flow. By analyzing historical data, the system establishes a baseline, identifying deviations as anomalies. The system calculates such by comparing current detections with historical data, flagging events that exceed two standard deviations from the mean. In our analysis, zero detections were excluded to avoid skewing the data.
Bird’s Eye View
@@
provides an accurate spatial representation, ideal for crowd management and area planning. The system normalizes object dimensions and uses a scale factor to accurately position objects within this view. Figure 5 compares the original and processed Bird’s Eye View from one camera, showing how this technique refines and clarifies spatial data.
Heat map
@@
visually represent data distribution, useful for monitoring congestion and spatial usage. The system generates 2D heatmaps from Bird’s Eye View coordinates, applying Gaussian smoothing for clarity. Figure 6 presents heatmaps for different 24-hour intervals. Graph (a) and (c) show weekday pedestrian movement, while graph (b) and (d) represent the weekend congestions at the same areas. Comparing the density pattern during weekday and weekend, reflect differences in area usage.
The detailed algorithms for all visualization tools are available in the supplementary materials for replication and further investigation.
Community Engagement
@@
To gauge public perception of such solution, we conducted seven rounds of survey studies in six public places across Charlotte, NC, during July and August 2023. A total of 410 respondents participated, providing insights into their concerns about safety and surveillance.
@@
The results indicated that respondents were most concerned about safety in parking lots (54%), public transit (44%), and entertainment venues (42%). Additionally, significant concerns were raised about potential biases and discrimination in current passive surveillance technologies, with these concerns being important or very important to respondents across all demographic groups, highlighting a widespread sensitivity to this issue.  [^1] This study has been previously published, but in adherence to the double-blind review policy of the AAAI conference, the reference will be provided in the camera-ready version upon acceptance.
@@
System Evaluation and Results
@@
(a) Latency Trends
(b) Throughput Trends
Figure 8: Latency and Throughput trends with respect to crowd densities across different nodes number running in parallelism.
Load Stress Evaluation
To rigorously assess the system’s performance under varying crowd densities and parallel node operations, we prepared ten videos with density levels ranging from 0 to 9 from CHAD dataset (Pazho et al. 2022). Each video lasted 150 seconds and operated at a frame rate of 30 frames per second. To ensure the statistical validity of our results, we averaged metrics over the central 100 batches in each run, excluding the initial 25 batches for system warm-up and the last 25 batches for cool-down effects.
Figure 8 graphically presents the trends for throughput and latency. The X-axis, labeled ”Density” (count), represents the average number of humans detected per batch (comprising 30 frames) for each input node number in the experiment. With one and four camera nodes, our system maintained latency under 10 seconds, showcasing stability regardless of scene crowdness. However, with eight nodes, latency increased beyond 20 seconds at higher density levels, particularly from level 5 onwards. A notable decrease in latency was observed with 12 nodes, especially at density levels 8 and 9, correlating with processing 108 individuals (12 nodes x 9 density).
Throughput consistently declined linearly with increasing density and node count, reaching a low of 4.56 FPS at the highest density level 9 with 12 nodes. The performance decrease is primarily due to the computational demands of pose estimator (Sun et al. 2019) processing high-density keypoints and is further constrained by CPU computation limits. As node count and density increase, these factors collectively create a bottleneck, leading to the observed decline in throughput and latency.
Real-World Endurance Evaluation
This evaluation assessed the system’s long-term stability using all 16 cameras in the testbed. We conducted extended trials with eight, twelve, and sixteen nodes running continuously for 21 hours and a week-long test using eight nodes to ensure the system’s robustness. This setup aimed to validate the system’s performance in realistic, extended operational conditions.
  • Twelve-Node Setup: Adds outdoor monitoring of parking lots.
As shown in Figure 7, the system’s performance over a week, running with eight cameras, demonstrated consistent latency and throughput results about crowd density. Specifically, the latency ranged from 2.5-7.8s, and throughput varied from 22.3-29.8 FPS. These findings confirm that irrespective of crowd density, the latency increased. At the same time, the FPS decreased, aligning with expectations and underscoring that the length of the videos does not significantly impact the system’s performance.
It is essential to note that the specific data spikes observed within this evaluation could stem from various factors, including network irregularities within the environment (as the input streams originate from IP cameras) or CPU memory usage nearing its limits. However, in light of the comprehensive experiment, these occasional data point spikes remain within acceptable bounds. The system consistently maintained optimal performance levels, showcasing its robustness and capacity to effectively manage and adapt to various operational scenarios.
For configurations with 8, 12, and 16 camera nodes, we observed similar trends over the 21-hour-long experiment with varying performance metrics: average latencies ranged from 2.6-4.8s, 5.3-6.5s, and 6.7-10.5s, respectively, while corresponding throughputs ranged from 28.5-26.5 FPS, 20.5-18 FPS, and 16.5-14.5 FPS. Detailed results with visualizations for 8, 12, and 16 camera nodes are provided in the supplementary materials for replication and further investigation.
(a) Object Anomaly
(b) Behavior Anomaly
Figure 9: Represent latency rug plot with different anomaly scenarios. 50 data points for each camera number
Physical-Cyber-Physical Evaluation (Anomaly detection)
This section evaluates the system’s Physical-Cyber-Physical (PCP) latency, representing the end-to-end time from when an anomaly appears to when the end user receives a notification. We assessed both object and behavioral anomaly detection (Markovitz et al. 2020b) using four, eight, twelve, and sixteen AI pipeline nodes reading input frames from the same camera node, conducting twenty experiments per anomaly type and collecting 400 data samples. The latency measures the time from detecting anomalies, such as high-priority objects or dangerous behaviors like fighting, to alerting the end user. A three-second transmission delay from cameras to AI Nodes was observed, slightly higher than other IP systems, due to our unique network setup. This delay is included in the overall PCP measurements, though manual recording of latency data may introduce minor discrepancies.
The results indicate that increasing the number of cameras significantly impacts PCP latency. For example, object anomaly latency increased from 4.7 seconds with four cameras to 19.68 seconds with sixteen, while behavioral anomaly latency rose from 9.34 seconds to 26.76 seconds. This increase is attributed to system traffic, multitasking overhead, and the processing capacity of the system.
Figure 9 displays the distribution of 400 data points from our PCP experiments. Graph (a) shows the distribution for object anomaly detection, with less variation in the four-camera setup compared to the sixteen-camera setup. Graph (b) illustrates the distribution for behavioral anomaly detection, where the data is more scattered with an increased number of cameras. The standard deviation for object detection latency increased from 0.65 to 1.12 seconds as the camera count increased, indicating a slight broadening of latency due to system complexity and resource contention.
This evaluation demonstrates the system’s capability to enhance safety and operational efficiency in real-world scenarios. The ability to notify end users within approximately 20 seconds is crucial for managing anomalies in public spaces, aiding in the prevention of potential dangers and improving resource allocation.

### v3

@@ 3.2 Server Module
Record Time Camera Class ID Bounding Box Feature Local ID Global ID Anomaly Score
@@
23:59:59 1 0 [x, y, w, h] Tensors 9999 1001 40
Server Module centralizes metadata and historical data from AI modules, stored in a MySQL database as outlined in Table 3. It handles global tracking and statistical analysis, using cosine similarity to re-identify individuals across cameras and analyze patterns [8]. The design prioritizes privacy by avoiding raw data transmission to the Cloud Module.
3.3 Cloud Module
@@
Cloud Module leverages cloud-native services for data storage, management, and API generation [36]. It minimizes lag between anomaly detection and notification, using a rule-based messaging service to send real-time alerts via email, text, or app notifications. A low-latency database supports real-time data access, while APIs generated through an application development kit optimize data retrieval [37, 4].
3.4 End User Devices
@@
End User Devices are designed to promptly notify users of detected anomalies via a smartphone application [9]. The app provides real-time data and analysis, ensuring consistent functionality across different devices and operating systems, thereby enhancing accessibility and delivering information efficiently.
4 Deployment and Setup
@@
In lab settings, each advanced software feature demonstrates significant potential individually; however, to fully realize its benefits, a comprehensive end-to-end evaluation in a real-world environment is essential. Such an evaluation allows for assessing the system’s load handling and endurance under actual conditions, ensuring that it can perform reliably and effectively when deployed, which might not be evident in a controlled environment.
Our study utilizes an existing surveillance network of 16 AXIS IP cameras, operating at 30 FPS with 720p resolution, across the campus, monitoring approximately 95,000 square feet of combined indoor and outdoor space. As shown in Fig. 3, thirteen indoor cameras are mounted and angled slightly downward to capture hallways, entryways, vending areas, and other common spaces, while three outdoor cameras, installed at 10 ft 8 in, oversee building exteriors, walkways, and surrounding grounds. All cameras feature varifocal lenses offering a 36–100° horizontal field of view, allowing flexible adjustments to accommodate wide-area observation. The system runs on a dedicated server with a 16-core CPU, 252 GB RAM, and four 24 GB VRAM GPUs.
The software components, encompassing the AI module and the Server Module as introduced in Section 3, are integrated into a dedicated server infrastructure. This server is interconnected within the local network of the testbed to enhance security measures. Video feeds from all cameras are transmitted utilizing RTSP. Data processed by each AI module is subsequently relayed to the database housed within the Server Module via TCP. The Cloud module server periodically aggregates data from the Server Module’s database for advanced data management purposes. The resultant outputs are then disseminated and updated across end-user applications
@@
As shown in Fig. 4, a sample visualized output from the AI module during a scenario with four operational cameras is presented, with privacy-preserving instance segmentation applied.
Figure 4: This represents an example of four pipelines running in the testbed, with IP cameras deployed at different locations.
5 Applications and Visualizations
This section illustrates how the data collected from the real world can be analyzed and visualized to enhance situational awareness in different civic applications. For example, visualizing pedestrian flow can inform the design of public spaces like intersection turnarounds to reduce congestion, while tracking real-time movement patterns helps prevent overcrowding and ensures safety in public events. These applications showcase how data-driven insights can lead to more efficient, safe, and well-planned communities.
5.1 Descriptive Data
@@
In our descriptive analysis, the Global ID from human feature data serves as the unique identifier for tracking across data streams. We present five key metrics to provide a general overview of traffic flow and occupancy trends: real-time count of people, hourly average number of people per camera per location, total number of people, and peak hour analysis over time. These metrics help monitor occupancy, understand distribution trends, and identify peak traffic times, aiding in resource allocation and emergency planning.
5.2 Situational Awareness
@@
Situational awareness is crucial for effective decision-making across various domains. Our study explores the four following key visualization techniques:
5.2.1 Occupancy Indicator
@@
The Occupancy Indicator provides insight into the number of people in a specific location by comparing current occupancy levels to historical data, which is particularly useful in emergencies or public health scenarios. To calculate occupancy, the system compares current detections with historical percentiles, categorizing occupancy as ”Low,” ”Normal,” or ”High.”
Figure 5: Comparing the Occupancy Indicator of the same camera during the same hour on two different days
. Graph (1) shows the Weekend. Graph (2) represents weekday.
Algorithm 2 describes a method for calculating and categorizing real-time occupancy levels across multiple camera locations. Operating on a continuously updated data frame (df), the algorithm processes each camera location sequentially. For each camera ID, it performs occupancy analysis at 5-second intervals throughout the specified record time. During each interval, the algorithm computes the current occupancy by determining the number of unique global IDs present in the data frame, effectively tracking the distinct individuals within each camera’s field of view.
Algorithm 2 Occupancy Indicator Algorithm
1:Input: data frame dfdf, historical_data
2:for each camera_id in dfdf do
3:  for every 5 seconds in record_time do
4:   current_#_of_people = length(unique(global_IDs))
5:  end for
6:  if current_#_of_people $\leq$ percentile(historical_data, 25) then
7:   occupancy = ”Low Occupancy”
8:  else if current_#_of_people $\leq$ percentile(historical_data, 75) then
9:   occupancy = ”Normal Occupancy”
10:  else
11:   occupancy = ”High Occupancy”
12:  end if
13:  update_historical_data(current_#_of_people)
14:end for
Occupancy levels are determined by comparing the current number of people against historical data percentiles. The algorithm classifies occupancy into three categories based on these thresholds: ’Low Occupancy’ when the count falls at or below the 25th percentile, ’Normal Occupancy’ when it is between the 25th and 75th percentiles, and ’High Occupancy’ when it exceeds the 75th percentile of historical data.
Fig. 5 compares occupancy indicators for the same camera during the same hour on two days. Graph 1 shows a weekend, with 2 people categorized as ”High Occupancy,” while Graph 2 shows the same number considered ”Normal” on a weekday.
Following each occupancy determination, the update historical data function incorporates the current number of people into the historical dataset. This continuous updating mechanism ensures the algorithm’s thresholds remain adaptive and relevant, reflecting recent occupancy patterns. The entire process executes across all camera IDs in the data frame, providing comprehensive occupancy monitoring across the entire surveillance network.
5.2.2 Statistical Anomaly
@@
Statistical Anomalies often indicate unusual traffic flow. By analyzing historical data, the system establishes a baseline, identifying deviations as anomalies. The system calculates such anomalies by comparing current detections with historical data, flagging events that exceed two standard deviations from the mean. In our analysis, zero detections were excluded to avoid skewing the data.
Algorithm 3 identifies statistically anomalous numbers of detected individuals for each camera relative to historical trends. Given that crowd densities vary across locations and times, these anomalies are calculated on an hourly basis. The approach involves constructing a normal distribution from historical hourly data, updated every 5 seconds as new detections occur. As outlined in Algorithm 3, the mean and standard deviation for each hour are calculated to characterize the data distribution. Current detections are then compared against this historical context. If the current number of detections exceeds two standard deviations from the mean, there is a less than 0.05 probability of such an occurrence, indicating with 95% confidence that it can be considered a statistical anomaly.
Algorithm 3 Statistical Anomaly Detection
1:Input: dfdf, start_timestart\_time, end_timeend\_time
2: Filter dfdf for camera ID in the specified time range
3: Initialize meanmean to 0 and stdstd to 1
4: Initialize detected_objectsdetected\_objects list
5:for each 5-second interval from start_timestart\_time to end_timeend\_time do
6:  Count unique detected objects in the interval
7:  if detected objects >>then
8:   Update detected_objectsdetected\_objects, meanmean, and stdstd
9:  end if
10:end for
We intentionally exclude detections of 0 in our analysis. The reason is that, within a 5-second window, the absence of detections (i.e., detecting no individuals) is more likely than detecting any individuals. Including 0 detections would skew the results toward zero, leading to lower thresholds for identifying anomalies and reducing the effectiveness of the statistical analysis.
5.2.3 Bird’s Eye View
@@
Bird’s Eye View provides an accurate spatial representation, ideal for crowd management and area planning. The system normalizes object dimensions and uses a scale factor to accurately position objects within this view.
Algorithm 4 describes the transformation of camera-detected objects into Bird’s Eye View coordinates, specifically designed for the AXIS P3225-VE Mk II Network Camera [^1] https://www.axis.com/products/axis-p3225-ve-mk-ii/support#technical-specifications. This transformation creates a top-down perspective of the surveillance area.
The transformation process begins with object dimension normalization, where width and height are divided by camera width and camera height, respectively. This step ensures dimensional independence from camera resolution. Subsequently, a scale factor is computed for each object based on its normalized height and the camera’s angular field of view parameters (min theta and max theta), which are obtained from the manufacturer’s specifications. This scaling mechanism preserves depth perception by appropriately adjusting object sizes based on their distance from the camera.
The algorithm then calculates each object’s centroid by averaging its width and height dimensions, establishing a precise reference point for positional mapping. The final Bird’s Eye View coordinates (BirdsEyeX and BirdsEyeY) are derived by applying the computed scale factor to these centroid coordinates, resulting in an accurate top-down spatial representation.
For temporal analysis, the algorithm implements a data filtering mechanism that processes records at 5-second intervals, tracking unique global IDs to determine object counts. This filtering ensures that analysis is performed only on relevant data within specified camera IDs and time frames.
Algorithm 4 Bird’s Eye View Transformation
1:Input: data frame dfdf, camera_width, camera_height, min_teta, max_teta
2: Compute normalized values: $df[^{\prime}normalized\_W^{\prime}]$, $df[^{\prime}normalized\_H^{\prime}]$
3:for each object in dfdf do
4:  Compute scale_factorscale\_factor using normalized_Hnormalized\_H
5:  Calculate centroid C_XC\_X and C_YC\_Y
6:  Compute BirdsEye_XBirdsEye\_X and BirdsEye_YBirdsEye\_Y using scale_factorscale\_factor and normalized values
7:end for
8:for camera_idcamera\_id from 1 to 8 do
9:  Define start_time and end_time
10:  Filter dfdf based on camera_idcamera\_id and time range
11:end for
Fig. 6 provides a comparative visualization between the original perspective and the processed Bird’s Eye View from one camera, captured simultaneously. In the initial view, each data point represents the XY coordinates of detected objects in the camera’s original field of view, showing approximately 12 distinct detections within a five-second interval. In the Bird’s Eye View, these XY coordinates are transformed, consolidating objects with the same global IDs into averaged positions. This results in nine distinct individuals, providing a top-down perspective that refines and simplifies the spatial arrangement.
((a))
((b))
Figure 6: Comparing different views from one camera. Graph (a) generated directly from the database shows the original camera view. Graph (b) represents the averaged bird’s eye view coordinates.
5.2.4 Heat map
@@
Heat maps visually represents data distribution, ann are useful for monitoring congestion and spatial usage. The system generates 2D heatmaps from Bird’s Eye View coordinates, applying Gaussian smoothing for clarity. Fig. 7 presents heatmaps for different 24-hour intervals. Graph (a) and (c) show weekday pedestrian movement, while Graph (b) and (d) represent the weekend congestion at the same areas. Comparing the density patterns during weekdays and weekends reflect differences in area usage.
((a))
((b))
((c))
((d))
Figure 7: Heatmap representations from two cameras during weekdays and weekends. Graph (a) and (c) represent the heatmaps during weekdays, while Graph (b) and (d) show congestion in the same areas during weekends.
Algorithm 5 is developed to generate heatmaps based on the data from cameras. We used the calculated Bird’s Eye View coordinates to generate heat maps. We processed data for each of the eight cameras, filtering records for every 24-hour window within the specified date range.
Algorithm 5 Heatmap Generation
1:Input: data frame dfdf, start_date, end_date
2: Compute date_range from start_date to end_date
3:for each date in date_range do
4:  for camera_idcamera\_id from 1 to 8 do
5:   Filter dfdf for current date and camera_idcamera\_id
6:   if filtered data is not empty then
7:    Determine max_xmax\_x and max_ymax\_y from filtered data
8:    Initialize heatmap array of size (max_y+1,max_x+1)(max\_y+1,max\_x+1)
9:    for each row in filtered data do
10:     Increment heatmap value at position (Bird’s Eye $Y$, Bird’s Eye $X$)
11:    end for
12:   end if
13:  end for
14:end for
The initial step involves defining a date range between a given start date and an end date. The algorithm then iterates over this date range and, for each date, processes data from each of the eight cameras. This systematic approach ensures comprehensive coverage, capturing the spatial dynamics of the monitored area over time.
For each date and camera combination, the dataset, $df$, is filtered to include only records that fall within the current date and are associated with the current camera. Once filtered, the algorithm checks if there’s any data available for processing.
If data is available for the given date and camera, the maximum Bird’s Eye View coordinates, max x and max y, are determined. These values define the dimensions of a 2D heatmap array, which is initialized with zeros. As the algorithm iterates over each row of the filtered data, the heatmap values are incremented based on the Bird’s Eye View positions of detected objects. Thus, by the end of this process, we obtain a series of heat maps, each capturing the spatial distribution of activity for a specific camera on a particular day, offering invaluable insights into the daily patterns of area usage and congestion. We applied Gaussian smoothing to the original heatmaps to enhance their clarity.
6 Community Engagement
@@
To gauge public perception of our solution, we conducted seven rounds of survey studies in six public places across Charlotte, NC, during July and August 2023. A total of 410 respondents participated, providing insights into their concerns about safety and surveillance.
@@
The results indicated that respondents were most concerned about safety in parking lots (54%), public transit (44%), and entertainment venues (42%). Additionally, significant concerns were raised about potential biases and discrimination in current passive surveillance technologies, being rated as important or very important to respondents across all demographic groups, highlighting a widespread sensitivity to this issue.  [^2] This study has been previously published, but in adherence to the double-blind review policy of the AAAI conference, the reference will be provided in the camera-ready version upon acceptance.
@@
7 System Evaluation and Results
7.1 Load Stress Evaluation
Our performance evaluation methodology employed a systematic testing approach using the CHAD dataset [38], comprising ten videos with crowd density levels ranging from 0 to 9. Each test video was 150 seconds in duration with a 30-frames-per-second capture rate. To ensure robust statistical analysis and minimize the impact of system initialization and termination effects, we focused our analysis on the central 100 batches of each test run, excluding the first 25 batches (warm-up period) and the final 25 batches (cool-down period) from our calculations.
((a))
((b))
Figure 8: Latency and Throughput trends with respect to crowd densities across different number of nodes running in parall. Graph (a) shows the latency trends, and Graph (b) shows the throughput trends.
Figure 8 illustrates the relationship between system performance metrics and crowd density. The X-axis represents the average number of individuals detected per 30-frame batch across varying input node configurations. System performance exhibited distinct patterns across different node configurations. With one to four camera nodes, the system maintained stable latency below 10 seconds across all density levels. However, the eight-node configurations showed significant latency increases exceeding 20 seconds at density levels above 5. Interestingly, the twelve-node configuration demonstrated improved latency at high density levels (8 and 9), particularly when processing approximately 108 individuals.
System throughput exhibited a linear decline correlating with increases in both density and node count, reaching its minimum of 4.56 FPS at density level 9 with 12 nodes. This performance degradation primarily stems from two factors: the computational intensity of the pose estimator [34] processing dense keypoint configurations, and CPU processing limitations. The combination of these factors creates a performance bottleneck that intensifies with increasing node count and crowd density.
7.2 Real-World Endurance Evaluation
This evaluation assessed the system’s long-term stability using all 16 cameras in the testbed. We conducted extended trials with eight, twelve, and sixteen nodes running continuously for 21 hours, and a week-long test using eight nodes to ensure the system’s robustness. This setup aimed to validate the system’s performance in realistic, extended operational conditions.
  • 1.
  • 2.
    Twelve-Node Setup: Adds outdoor monitoring of parking lots.
  • 3.
Figures 9, 10, and 11 present comprehensive analyses of system latency and throughput patterns across varying crowd densities during a 21-hour operational period. The evaluation, conducted from 2:00 p.m. to 11:00 a.m. the following day, was tested with three camera node configurations. Our observations revealed a notable decrease in system latency when the server node’s database approached 50,000 queries. To maintain optimal performance and ensure data privacy, we implemented an automatic database reset mechanism that activates either upon reaching 50,000 queries or after 24 hours of continuous operation.
Figure 9: Latency and throughput trends concerning crowd densities during a 21-hour-length period for 8 camera nodes
Figure 10: Latency and throughput trends concerning crowd densities during a 21-hour-length period for 12 camera nodes
Figure 11: Latency and throughput trends concerning crowd densities during a 21-hour-length period for 16 camera nodes
The data presented in Figures 9, 10, and 11 represent values averaged over 60-second intervals. The ’Density’ metric reflects the aggregate number of human detections per camera node during these one-minute sampling windows. The relatively low human traffic observed, particularly from indoor camera nodes, can be attributed to the data collection period coinciding with the summer break period.
@@ 7.2 Real-World Endurance Evaluation
Figure 12 illustrates the system’s week-long performance using an eight-camera configuration. The results demonstrate consistent correlations between crowd density and performance metrics, with latency ranging from 2.5-7.8 seconds and throughput varying between 22.3-29.8 FPS. These findings confirm our expected inverse relationship between crowd density and system performance, while demonstrating that video duration has minimal impact on overall system performance.
Within the dataset, occasional performance spikes were observed, potentially attributable to environmental network fluctuations from IP camera streams or periods of peak CPU memory utilization. However, these variations remained within acceptable operational parameters, demonstrating the system’s resilience under diverse conditions. The extended 21-hour evaluation across different camera configurations revealed consistent performance patterns:
  • 1.
    8 nodes: latency of 2.6-4.8s with throughput of 28.5-26.5 FPS
  • 2.
    12 nodes: latency of 5.3-6.5s with throughput of 20.5-18 FPS
  • 3.
    16 nodes: latency of 6.7-10.5s with throughput of 16.5-14.5 FPS
The system demonstrated consistent performance under diverse operational conditions, validating its robustness and adaptability.
7.3 Physical-Cyber-Physical Evaluation (Anomaly detection)
This section evaluates the system’s Physical-Cyber-Physical (PCP) latency, representing the end-to-end time from when an anomaly appears to when the end user receives a notification. To showcase the versatility and robustness of our AI-driven surveillance system, we established Testbed B with 12 CCTV cameras—3 outdoor and 9 indoor—at a busy market shop, featuring a 32-core CPU, 512 GB RAM, and two 48 GB VRAM GPUs. This setup complements Testbed A, providing diverse environmental conditions and hardware configurations for testing anomaly detection. For a comparative baseline, we deployed the same pipeline on AWS cloud, using an EC2 p2.xlarge instance with a single GPU to process video streams from IoT IP cameras over Real-Time Streaming Protocol. Metadata is stored in Amazon RDS, queried via Athena, and notifications are sent through SNS, mirroring the edge architecture for a direct performance comparison. Fig. 13 illustrates example camera views from two testbeds without any human presence.
By comparing the edge-based setups with the cloud deployment, we specifically examined end-to-end latency differences. In the cloud setup, the EC2 instance consolidates local and global processing nodes, with Amazon services enabling real-time analytics and alerts. This configuration highlights how offloading inference to the cloud impacts response times, with results from the p2.xlarge instance serving as a reference point. While alternative AWS configurations could yield different latencies and costs, this setup provides valuable insights into the trade-offs of cloud processing, aiding in the optimization of surveillance system design based on operational needs.
Figure 13: Example footages of the cameras in different locations
We assessed both object and behavioral anomaly detection [29] using four, eight, twelve, and sixteen AI pipeline nodes reading input frames from the same camera node, conducting fifty experiments per anomaly type and collecting 700 data samples.
PCP latency measures the time from detecting anomalies, such as high-priority objects or dangerous behaviors like fighting, to alerting the end user. As mentioned in Sec 3, the object detector can detect object-of-interest classes as object anomaly and behavior anomaly detector is trained on normal movement sequences and evaluates incoming sequences to see whether the movement is normal or not. For evaluation, we used predefined test cases: prohibited items were placed in predetermined locations, and staged physical confrontations were enacted by trained, authorized participants within the detection range of the camera under examination. These controlled scenarios ensured consistent triggering of the system’s anomaly detection mechanisms, enabling accurate measurement of end-to-end alert latency.
Various strategies were implemented to reflect the system’s versatility across the two testbeds. For testbed A, although the number of camera nodes was increased, multiple streams were duplicated from the same camera, ensuring that crowd density remained controlled. By contrast, testbed B, employed different camera streams onsite when increasing the node count, resulting in an uncontrolled number of individuals entering the field of view, thereby simulating real-world conditions.
About three-second transmission delay from cameras to AI nodes was observed for testbed A, slightly higher than other IP systems, due to its unique network setup. This delay is included in the overall PCP measurements, though manual recording of latency data may introduce minor discrepancies.
Table 4 provides more detailed insight about P-C-P evaluation results. This evaluation demonstrates the system’s capability to enhance safety and operational efficiency in real-world scenarios. The ability to notify end users within approximately 20 seconds enables transforming the existing CCTV systems from passive video cameras to proactive systems that can notify the stakeholders about the events that are happening.
The results indicate that increasing the number of cameras significantly impacts PCP latency. For example, object anomaly latency increased from 4.7 seconds with four cameras to 19.68 seconds with sixteen, while behavioral anomaly latency rose from 9.34 seconds to 26.76 seconds in testbed A. This increase is attributed to higher system traffic, multitasking overhead, and the overall processing capacity. Tests at Testbed B confirmed these trends, illustrating how a more dynamic environment and increased workload can further affect end-to-end alert times. In contrast, the cloud-based system offers significant advantages in terms of performance stability. Unlike the edge system, which is constrained by its local processing capabilities and experiences substantial latency increases as more cameras are added, the cloud infrastructure can dynamically allocate additional processing instances to handle the increased workload from more cameras. However, the result of 21.38 seconds for object anomaly detection and 26.5 seconds for behavior anomaly detection is not competitive with the edge based system in the experiment.
Table 4: Statistical PCP latency data for object and behavior anomaly for different number of cameras in different environment
Location A (Controlled) Location B (Open) Full Cloud
Nodes Mean (s) Min (s) Max (s)
Standard
deviation
Mean (s) Min (s) Max (s)
Standard
deviation
Mean(s)
Object Anomaly 4 4.7 3.06 6.4 0.55 3.372 2.29 5.01 0.64 21.38
8 10.99 9.18 12.55 0.85 8.924 7.46 10.61 0.71
12 15.76 14.05 18.31 1.05 13.037 10.28 16.93 1.27
16 19.68 17.67 21.6 1.13 - - - -
Behavior Anomaly 4 9.34 7.81 10.41 0.65 6.09 4.78 7.37 0.6 26.5
8 14.53 13.23 16.46 0.71 11.62 10.75 13.3 0.62
12 18.45 16.32 20.35 0.97 17.68 15.42 20.55 1.06
16 26.76 24.23 29.54 1.52 - - - -